# QuantFace: Low-Bit Post-Training Quantization for One-Step Diffusion Face Restoration

### v1

@@ 4.2 Main Results
Table 4: Ablation studies on different components. Experiments are conducted on the synthetic dataset CelebA-Test [57] and the bit setting is W4A4. The best and the second best results are colored red and blue, respectively. Rotation stands for rotation-scaling channel balancing, and MP stands for mixed-precision configuration generated by adaptive bit-width allocation.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Wider 0175 OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption WebPhoto 00064 OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption LFW A. Kwasniewski OSDFace [12] MinMax [38] Q-Diffusion [17] EfficientDM [21] PassionSR [14] SVDQuant [15] QuantFace (ours)
Figure 7: Visual performance comparison on real-world datasets. Please zoom in for a better view. Our QuantFace is under bit-width W4A3.98 and other methods are under W4A4.
@@ 4.3 Ablation Studies
Rotaion-Scaling Channel Balancing.
As illustrated in Fig. 4, solely applying scaling leads to uneven data distribution across channels, making it challenging to preserve the original data distribution after quantization. On the other hand, using the Randomized Hadamard Transform (RHT) alone fails to eliminate extreme outliers in the activations. By combining both techniques, we can effectively maintain the data distribution while reducing quantization errors. Consequently, we are able to enhance the final face restoration performance.
@@
The results presented in the Tab. 4 demonstrate that QD-LoRA effectively maintains a balance between quantization and distillation. QD-LoRA neither struggles to align with the FP model nor overfits to the limited calibration dataset. Therefore, models can generate excellent human face images after quantization. Results in Tab. 4 also reveal that QD-LoRA surpasses direct optimization on the SVD initialized low-rank branch.
@@
Various results shown in Tab. 1 and Tab. 4 demonstrate that our mixed-precision configuration enhances reference-based metrics on the synthetic dataset. The results indicate that the generated face images align closely with real-world face images. Visually, our configuration significantly improves the detail of generated facial features, including eyes and teeth.
@@ 5 Conclusion
We introduce QuantFace, an effective quantization framework for one-step diffusion face restoration models. QuantFace not only reduces quantization error but also facilitates the distillation from the full-precision model to the quantized model. Moreover, QuantFace promotes reasonable resource allocation. Extensive experiments demonstrate that QuantFace outperforms existing SOTA methods across a variety of metrics. By maintaining a balance between computational overhead and restoration quality, QuantFace makes the efficient deployment of advanced face restoration models possible.

### v2

@@ 4.2 Main Results
SVDQuant 0.5192 0.4143 76.3504 10.387 84.91
+ GAN Loss 0.5374 0.4249 76.8697 8.852 80.71
+ QD-LoRA 0.6321 0.4060 77.3236 9.344 58.88
+ Rotation 0.6916 0.3709 69.2888 6.195 58.34
+ Mixed-Precision 0.6984 0.3518 65.6640 5.927 53.18
Table 4: Ablation studies on different components starting from SVDQuant-W4A4. Experiments are conducted on the synthetic dataset CelebA-Test [CelebA-Test].
@@ 4.3 Ablation Studies
Rotation-Scaling Channel Balancing.
As illustrated in Fig. 4, solely applying scaling leads to uneven data distribution across channels, making it challenging to preserve the original data distribution after quantization. On the other hand, using the RHT alone fails to eliminate extreme outliers in the activations. By combining both techniques, we can effectively maintain the data distribution while reducing quantization errors. Consequently, we are able to enhance the final face restoration performance.
@@
The results presented in the Tab. 4 demonstrate that QD-LoRA effectively maintains a balance between quantization and distillation. QD-LoRA neither struggles to align with the FP model nor overfits to the limited calibration dataset. Therefore, models can generate excellent human face images after quantization. Results in Tab. 4 also reveal that QD-LoRA surpasses direct optimization on the SVD low-rank branch.
@@
Various results shown in Tab. 1 and Tab. 4 demonstrate that our mixed-precision configuration enhances reference-based metrics on the synthetic dataset. The results indicate that the generated face images align closely with real-world face images. Visually, our configuration significantly improves the detail of the generated facial features, including eyes and teeth.
@@ 5 Conclusion
We introduce QuantFace, an effective quantization framework for one‑step diffusion face restoration models. QuantFace not only reduces quantization error but also facilitates the distillation from the full-precision model to the quantized model. Moreover, QuantFace promotes reasonable resource allocation. Extensive experiments demonstrate that QuantFace outperforms existing SOTA methods across a variety of metrics. By maintaining a balance between computational overhead and restoration quality, QuantFace makes the efficient deployment of advanced face restoration models possible.